OpenStudy (anonymous):

There are 28 students in Mrs. Stanley's class. 1/7 of them got a "B" in math class. How many students got a "B"? Answer in simplest form. Plz Help!! Medals will be given & I'll follow yuh [: Kay So Please! :}

OpenStudy (anonymous):

28*1/7=4 students go a B

OpenStudy (anonymous):



OpenStudy (anonymous):

28*(1/7)=28/7=4

OpenStudy (anonymous):

Thanks Guys [[; Gonna follow! Plz follow bak {;

OpenStudy (anonymous):

*I mean be yur fan. Be my fans plz [[: